Change 5.21E+0 kg of silver to atoms of silver. Respond with the correct number of significant figures in scientific notation (Use E notation and only 1 digit before decimal e.g. 2.5E5 for 2.5 x 10⁵)
Answer #1

Molar mass of Ag = 107.9 g/mol

mass(Ag)= 5.21 Kg

= 5210.0 g

use:

number of mol of Ag,

n = mass of Ag/molar mass of Ag

=(5.21*10^3 g)/(1.079*10^2 g/mol)

= 48.29 mol

use:

number of atoms = number of moles * Avogadro’s number

= 48.29 * 6.022*10^23 atoms

= 2.908*10^25 atoms

Answer: 2.91E25